OpenStudy (anonymous):

if 7a=3 and 3b=7, what is the value a/b

OpenStudy (anonymous):

9/49

OpenStudy (gorv):

7a=3 means a=3/7 3b=7 means b=7/3 \[\frac{ a }{ b }=\frac{ \frac{ 3 }{ 7 } }{ \frac{ 7 }{ 3 } }=\frac{ 3 }{ 7 }*\frac{ 3 }{ 7}\]

OpenStudy (gorv):

a/b=9/49